A piece of equipment has a lifespan of T which is exponentially distributed with mean 20. A warranty pays (20 – T) if the failure occurs prior to time 20 and 0 otherwise. Find the variance of the warranty payment.

  1. 52

  2. 75

  3. 150

  4. 199

  5. 331
Answer #1

let x be payment

here E(X) = (20-t)f(t) dt = (20-t)*(1/20)*e-t/20 dt =7.3576

E(X2)= (20-t)2f(t) dt = (20-t)2*(1/20)*e-t/20 dt =105.6964

Var(X)=E(X2)-(E(X))2 =105.6964-7.3576^2 =52